End of Year $ 13,000 13,000 13,000 13,000 13,000 $ 13,000 4 $13,000 13,000 13,000 13,000 13,000 13,000 65,000 10 13,000 Assuming an annual discount rate of 18 percent, find the present value of each investment.
Answer #1

Present value of cash flows=cash flows*Present value of discounting factor(rate%,time period)

A:

Present value=13000/1.18+13000/1.18^2+13000/1.18^3+13000/1.18^4+13000/1.18^5

=13000[1/1.18+1/1.18^2+.............+1/1.18^5]

=$13000*3.127171021

=$40653.22(Approx).

B:

Present value=13000/1.18^5+13000/1.18^6+.............+13000/1.18^10

=13000[1/1.18^5+1/1.18^6+...............+1/1.18^10]

=($13000*1.80402449)

which is equal to

=$23452.32(Approx).

C:

Present value=13000/1.18+65000/1.18^6+13000/1.18^10

=$37578.84(Approx).